there are 3 red pens, 4 blue pens and 5 black pens in a box, Sameena picks a pen at random from the box. What is the probability that the pen is NOT black


1
Expert's answer
2021-07-06T13:55:01-0400

Total number of pens = 3+4+5 = 12

The number of NOT black pens = 3+4=7

P(pen is NOT black) "= \frac{7}{12}"